OpenStudy (anonymous):

Examine the following unbalanced equation. HCl (g) + O2 Cl2 (g) + H2O (g) Assuming that all the gases in this reaction are at STP, what number of molecules of hydrogen chloride gas are needed to completely react with oxygen to yield 50.00 mL of chlorine gas?

sam (.sam.):

First you need to balance the equation

sam (.sam.):

\[\begin{array}{cc} 4 & \text{H}\text{Cl} \\\end{array}\text{ + }\begin{array}{c} \text{O}_2 \\\end{array}\text{ $\longrightarrow $ }\begin{array}{cc} 2 & \text{Cl}_2 \\\end{array}\text{ + }\begin{array}{cc} 2 & \text{H}_2\text{O} \\\end{array}\]

sam (.sam.):

\[50ml=0.050L=0.05dm^3\]

sam (.sam.):

\[0.05dm^3Cl_2\times \frac{1molCl_2}{24dm^3Cl_2}\times\frac{4molHCl}{2molCl_2}\times \frac{2(6.02\times10^{23})moleculesHCl}{1molHCl}=?\]

OpenStudy (anonymous):

is the answer 599*10^24 molecules?